Two particles, A and B, are in uniform circular mo
Answer #1

a[A] = r[A]w[A]^2
a[B] = r[B]w[B]^2
where w[A] and w[B] are the angular velocities of A and B.
a[A] = 4.7*a[B]

T[A] = 2*pi/w[A]
T[B] = 2*pi/w[B] = 2.8*T[A]
T[B]/T[A] = 2.8 = w[A]/w[B]

w[A]^2/w[B]^2 = 7.84
a[A]/a[B] = 4.7 = r[A]w[A]^2/r[B]w[B]^2 = 7.84(r[A]/r[B])
r[A]/r[B] = 4.7/7.84 = 0.599